OpenStudy (anonymous):

Two cars leave Denver at the same time and travel in opposite direction. one car travels 5mi/h faster that the other car. The care are 575 mi apart in 5 hours. How fast is each car traveling? 2nd attempt for this question...

OpenStudy (saifoo.khan):

So we got it wrong last time? :o

OpenStudy (anonymous):

the answer i got was not any of the options for the multiple choices

OpenStudy (saifoo.khan):

Oh. let me try..

OpenStudy (anonymous):

thx

OpenStudy (saifoo.khan):

@yummydum did it wrong.

OpenStudy (anonymous):

i got it...gimme a sec

OpenStudy (saifoo.khan):

He didn't used the total time in his solution.

OpenStudy (anonymous):

what formula did u come up with?

OpenStudy (anonymous):

\[d_1=5r\]\[d_2=5(r+5)\implies d_2=5r+25\]\[d_1+d_2=575\] \[5r+5r+25=575\]\[10r=550\]\[r=55\] \[d_1=55mph\]\[d_2=60mph\]

OpenStudy (anonymous):

@justinsmom better?

OpenStudy (anonymous):

ok thats one of the options... hope he can figure other two out with info i got earlier... thanks again

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

I don't get it @yummydum , explain.

OpenStudy (anonymous):

kyun bhai??

OpenStudy (saifoo.khan):

5r+25?

OpenStudy (anonymous):

\[5(r+5)\implies 5r+25\]

OpenStudy (anonymous):

since the second car is traveling 5mph faster than the first

OpenStudy (saifoo.khan):

why multiply it by 5?

OpenStudy (anonymous):

because that is the value of t(time)....t=5hrs

OpenStudy (saifoo.khan):

So?

OpenStudy (anonymous):

the formula is d=rt d=r5

OpenStudy (anonymous):

that is for the first car and the second car is 5 plus the rate so that makes it d=(r+5)5

OpenStudy (saifoo.khan):

how about? \[Speed = \frac{distance}{time}\]\[r+r+5 = \frac{575}{5}\]

OpenStudy (saifoo.khan):

Ah, same thing. Mine's easier..

OpenStudy (anonymous):

yea thats the same thing 2r+5=115 2r=110 r=55

OpenStudy (saifoo.khan):

Still, mine's easier.

OpenStudy (anonymous):

yes..it is...khush?

OpenStudy (saifoo.khan):

``` yummydummy

OpenStudy (anonymous):

o.O

OpenStudy (saifoo.khan):

``` heck yeah! :D

OpenStudy (anonymous):

umm...theek tou ho...aik dum kya hua? :P

OpenStudy (saifoo.khan):

Lol, kuch nahi hua bhai. Sab waisa he hai.

OpenStudy (anonymous):

acha